# PEFT-Bench: A Parameter-Efficient Fine-Tuning Methods Benchmark

### v1

@@ Abstract
Despite the state-of-the-art performance of Large Language Models (LLMs) achieved on many tasks, their massive scale often leads to high computational and environmental costs, limiting their accessibility. Parameter-efficient fine-tuning (PEFT) methods address this challenge by reducing the number of trainable parameters while maintaining strong downstream performance. Despite the increased development in PEFT methods, current evaluations remain limited (in terms of evaluated models and datasets) and difficult to reproduce. To bridge this gap, we introduce PEFT-Bench, a unified end-to-end benchmark for evaluating diverse PEFT methods on autoregressive LLMs. We demonstrate its usage across 27 NLP datasets and 6 PEFT methods. To account for different PEFT training and inference factors, we also introduce the PEFT Soft Score Penalties (PSCP) metric, which takes trainable parameters, inference speed, and training memory usage into account.
@@ 1 Introduction
Large Language Models (LLMs) (Radford et al., 2019; Achiam et al., 2023; Dubey et al., 2024; Jiang et al., 2023; Team et al., 2024) achieve remarkable results in many Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks. This significant increase in performance was possible due to the introduction of the attention mechanism forming the transformer architecture (Vaswani, 2017). To achieve such performance, many currently best-performing models contain a vast number of attention blocks, increasing the number of trainable parameters to millions (Devlin et al., 2019), billions (Radford et al., 2019), or even trillions (Fedus et al., 2022). With the increase in parameters, the computational and data requirements for training also increased, making the development of LLMs unavailable for many academic institutions or practitioners, as well as producing a bigger carbon footprint. To address this issue, a new paradigm was introduced, denoted as “pre-train, fine-tune”, in which the language model is at first pre-trained on large amounts of raw data and then, in a fully supervised way, adapted to the downstream task (Liu et al., 2023a). This standard fine-tuning still requires substantial amounts of resources and, furthermore, it requires significant quantities of data and storage to store fine-tuned weights for each task.
@@
To ultimately improve and unify the evaluation of PEFT methods, we propose a new benchmark PEFT-Bench, which bridges the gap in the comparison of PEFT methods. We simultaneously introduce the PEFT-Factory framework [^1] To promote the replicability and usage of our benchmark, we provide the repository to our PEFT-Factory framework., which provides a necessary underlying technological support for execution of the PEFT-Bench benchmark.
@@
  • We introduce the PEFT-Bench, an end-to-end benchmark that defines the datasets, metrics, and methodology for evaluating PEFT methods in NLP in a fair and consistent environment. To the best of our knowledge, we are the first to provide such a benchmark in NLU and NLG tasks for PEFT methods with autoregressive LLMs. To demonstrate its use, we evaluate 6 different PEFT methods in terms of efficiency and stability with limited data on 27 NLP datasets.
@@
  • Finally, we propose the PEFT Soft Cost Penalties metric (PSCP), which introduces a number of trainable parameters, memory usage, and inference speed in the final score calculation. PSCP reflects the practical feasibility of deploying PEFT methods in real-world scenarios, where efficiency and resource constraints are often as critical as task performance.
@@ 2 Related Work
Many new PEFT methods are primarily evaluated on GLUE and SuperGLUE benchmarks and often include the main results only on non-autoregressive models (Tang et al., 2025; Zhang et al., 2025b, a) (e.g., T5 (Raffel et al., 2020), RoBERTa (Liu et al., 2019)) or include only limited evaluation on autoregressive models Lan et al. (2025); Shi and Lipani (2024); Belanec et al. (2025). Although some of the new parametrization-based methods (e.g., SVFT (Lingam et al., 2024), SVF (Sun et al., 2025)) are already evaluated on autoregressive models and on natural language generation tasks, a more unified and overarching evaluation is still missing.
@@
Additionally, due to low baseline replicability, we have been witnessing a trend in new soft prompt-based methods to draw numerical results directly from the related works (without actually rerunning the experiments) (Zhang et al., 2025b; Tang et al., 2025). This practice is often not only methodologically incorrect (if the authors do not perfectly match the original environment used to evaluate baselines), but also weakens the results of the proposed method.
@@ 3 PEFT-Bench
Our PEFT-Bench is constructed out of multiple steps, which are visualised by the diagram in the Figure 1. At a high level, it consists of three main components: 1) a diverse collection of datasets and tasks; 2) representative language models together with a set of PEFT methods; and 3) a set of evaluation metrics.
@@ 3.1.1 Natural Language Understanding and Reasoning Tasks
PEFT-Bench includes 8 datasets from GLUE, separated into 6 tasks, namely natural language inference (NLI)MNLI (Williams et al., 2018), QNLI (Rajpurkar et al., 2016), RTE (Dagan et al., 2005; Bar Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009); paraphrase classificationQQP  [^2] https://quoradata.quora.com/First-Quora-Dataset-Release-Question-Pairs, MRPC (Dolan and Brockett, 2005); sentiment classificationSST-2 (Socher et al., 2013); sentence similaritySTS-B (Cer et al., 2017) and acceptability classificationCoLA (Warstadt et al., 2019).
@@
Additionally, PEFT-Bench also includes 7 datasets from SuperGLUE, separated into 4 tasks, namely natural language inference (NLI)CB (De Marneffe et al., 2019); question answeringMultiRC (Khashabi et al., 2018), ReCoRD (Zhang et al., 2018), BoolQ (Clark et al., 2019), COPA (Roemmele et al., 2011); word sense disambiguationWiC (Pilehvar and Camacho-Collados, 2019) and coreference resolutionWSC (Levesque et al., 2011);
@@
PEFT-Bench also includes datasets that are, to the best of our knowledge, not part of any well-known benchmarks or are benchmarks on its own, but are still popular and commonly used to evaluate NLU and reasoning in LLMs, totaling 6 datasets separated into 3 tasks, namely question answeringMMLU (Hendrycks et al., 2021), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), OBQA (Khot et al., 2019); natural language inference (NLI)HellaSwag (Zellers et al., 2019); commonsense reasoningWinoGrande (Sakaguchi et al., 2021);
@@ 3.1.2 Math Tasks
Another popular task that is used for LLM evaluation is mathematical problem-solving. The difficulty of datasets for such evaluation can range from simple question answering to full mathematical reasoning and problem-solving. PEFT-Bench includes 3 datasets for mathematical problem solving separated into 3 tasks, namely question answeringMathQA (Amini et al., 2019); math word problemsGSM8K (Cobbe et al., 2021) and simple math problemsSVAMP (Patel et al., 2021).
@@ 3.2 PEFT Methods and Pretrained Models
With the popularization of PEFT, many new methods are being introduced within a short period of time (Xu et al., 2023; Prottasha et al., 2025). Therefore, it is computationally expensive to evaluate them all. We design our PEFT method selection in PEFT-Bench to cover additive PEFT, reparametrized PEFT, and selective PEFT categories (the categorization of methods that we have selected can be found in Figure 3). To account for each category, we select IA3 (Liu et al., 2022a), Prompt Tuning (Lester et al., 2021), P-Tuning (Liu et al., 2022b), and Prefix Tuning (Li and Liang, 2021) from additive PEFT methods, where IA3 is an adapter-based method, and the rest are soft prompt-based methods. Consequently, we select LoRA (Hu et al., 2022) from reparametrized PEFT methods and LNTuning (Zhao et al., 2024) from selective PEFT methods. These methods are widely adopted, frequently used as baselines, and together represent one of the most influential and diverse approaches within each PEFT category. We provide a more detailed description of each selected PEFT method in Appendix C.
@@ 3.3 Evaluation metrics
Current works on PEFT methods often measure solely the accuracy performance or the number of trainable parameters, which makes it harder to compare. Moreover, related works only rarely report the memory and inference efficiency in their results. These are the key factors that may affect the preference for using a specific method over another within a certain computational budget. At the same time, for efficient comparison of benchmarked methods, we need to aggregate performance and such additional PEFT training and inference factors into a single metric. To this end, we propose a new PEFT-bencharking metric called PEFT Soft Score Penalties (PSCP).
Within this metric, we propose to penalize the performance when the costs behind a certain factor are too high. Therefore, we need to scale the performance on the specific task $P_{t}$ with a cost for that factor, similarly to the Cobb-Douglass production function (Cobb and Douglas, 1928) for $n$ costs:
@@
where $\beta$ is the importance of each cost, which can also be viewed as a corresponding weight of each penalty.
In our case, we consider the following factors: 1) the number of trainable parameters $M_{p}\in\mathbb{N}$; 2) inference speed $M_{f}\in\mathbb{R}^{+}$ (i.e., floating point operations); and 3) average memory usage during training $M_{m}\in\mathbb{R}^{+}$. Since, these factors are not normalized and cannot be used as penalties on their own, we normalize $M_{p}$, $M_{f}$, $M_{m}$ using reference constants $\frac{M_{p}}{C_{p}}$, $\frac{M_{f}}{C_{f}}$, $\frac{M_{m}}{C_{m}}$, where $C_{p}\in\mathbb{N}$, $C_{f}\in\mathbb{R}^{+}$ and $C_{f}\in\mathbb{R}^{+}$. The constants can be set either to a large value that we would not like to achieve (e.g., number of parameters close to full model training, double the inference floating point operations, or maximum memory available for training), or alternatively, they can also be set to the median or geometric mean of all available $M$ values from a specific factor. If we plug this normalization into Equation 1, we would end up with the intermediate equation 2.
@@
This, however, would rescale the performance $P_{t}$ to infinity if $M>>C$, therefore bound it to the interval $[0,1)$ with the reciprocal function (i.e., $x^{-}1$). For this reason, we also need to add 1, so that the result of the reciprocal does not go to infinity when $M<<C$. The final form of our PEFT Soft Cost Penalties (PSCP) metric can be found in the Equation 3.
@@
The proposed metric is easily extensible if additional factors are of interest. In our case, we consider factors that we want to minimize (i.e., the number of parameters, inference speed, and training memory). If we want to include different factors that we want to maximize (i.e., higher is better), we need to switch the numerator and denominator of the fraction for that particular cost calculation.
In addition, the PSCP allows setting the importance of each factor by setting the $\beta$ value to different values, which makes the PSCP metric further configurable and adaptable for different use cases. Depending on which efficiency factor we want to put emphasis on, we can set the importance $\beta$ to a value between 0 and 1. While it will keep the penalty still in the interval $[0,1)$, the resulting penalty value will be closer to 1, which finally decreases the influence of the penalty on the measured performance $P_{t}$. In addition, we can set the importance $\beta$ of specific factors to a number greater than 1, which will make the penalty go closer to zero, and therefore, more influential for that factor. This specific behaviour can be seen in the Appendix B in Table 9(a), where we decrease the importance of the number of trainable parameters, and therefore, decrease the penalty, which made LoRA a method with the highest PSCP score. Alternatively, if the importance factors are set equally, only the magnitude of the PSCP changes, not the order of methods. This can be seen in Appendix B in Table 10.
@@ 3.4 PEFT-Factory
To increase the replicability, modularity, and usability of the PEFT-Bench, we simultaneously propose PEFT-Factory, a framework for efficient training of autoregressive LLMs, based on one of the popular LLaMa-Factory LLM training tools (Zheng et al., 2024). PEFT-Factory includes new off-the-shelf PEFT methods from the HuggingFace PEFT library (Mangrulkar et al., 2022) and also provides an easy-to-use interface to add newly created PEFT methods to our framework. In addition, PEFT-Factory also includes all PEFT-Bench datasets, from which some are specifically adapted for autoregressive generation and classification (e.g., GLUE and SuperGLUE). In addition, PEFT-Factory also implements the PSCP metric together with standard metrics and provides automated evaluation and results comparison.
Consequently, PEFT-Factory can be viewed as an instrument of PEFT-Bench that automates the training of different methods in a simple and configurable way. In this manner, PEFT-Factory takes just two configuration files (one for training and one for evaluation) as input, and returns fine-tuned adapter weights alongside the performance metrics. This specific property of PEFT-Factory allows an identical experimental environment during multiple runs (when equivalent hardware configurations are utilized). For this reason, we create a specific training and evaluation configuration for each model and method combination during the PEFT-Bench execution.
@@ 4 Experimental Setup
In all datasets, we utilize 10% randomly selected from the training set as a validation set and the original validation set as a test set for evaluation. During the validation phase, we measure validation loss and save the weights of the best validation loss for every 5% of the training steps. We train for 10 epochs with a batch size of 4. We utilize the AdamW (Loshchilov and Hutter, 2019) optimizer with a 0.1 warmup ratio, a weight decay of 1.0e-5, a learning rate of 5.0e-5, and a cosine learning rate scheduler. If not specified otherwise, we keep the training-related hyperparameters equal for all of the experiments and change only the PEFT method hyperparameters.
For IA3, we apply adapters for key, value, and downsample feedforward modules of the transformer architecture. For Prompt Tuning, we fine-tune the soft prompt with a length of 100 tokens and initialize it randomly from the vocabulary of the trained model. For P-Tuning, we also fine-tune the soft prompt with a length of 100 tokens, but we reparametrize the soft prompts with an LSTM model with a hidden size of 768. For Prefix Tuning, we fine-tune the soft prompt with a length of 32 tokens, and we project the soft prompt with a small MLP with a hidden size of 512. For LoRA hyperparameters, we set the rank to 16, alpha to 16, and dropout to 0.05. Similarly to IA3, we also reparametrize the key, value, and downsample matrices of the transformer architecture. For LNTuning, we leave all of the settings to their default values (the modules are selected based on the model architecture). For the stability experiments, we rerun training for 5 different random seeds, and we train each method for 20 epochs on each dataset.
To calculate the PSCP metric, we calculate $P_{t}$ as an average of the performance metric (depending on the particular dataset) across all 5 runs. As all the performance metrics for each dataset have a scale of <0,1><0,1>. This allows us to average all individual performance scores into a single value $P_{t}$ in a fair manner (as it is already done in the established benchmarks as well, such as GLUE and SuperGLUE). We set $\beta$ for all factors to 1. For the efficiency factors, we set $M_{p}$ as the number of trainable parameters of the PEFT model (counting the parameters that are enabled in gradient descent). Consequently, calculate $M_{f}$ as the difference between base model FLOPs and PEFT model FLOPs. Lastly, we also calculate the $M_{m}$ as the average maximum memory across all runs.
Moreover, we set all reference constants to values close to full fine-tuning of the base model, in which case it would be more beneficial to fine-tune the full model instead of the PEFT model. Specifically, we set $C_{p}=5\times 10^{8}$, as some of the smallest LLMs contain 0.5 billion parameters (Yang et al., 2024), $C_{f}=10$, as a difference in 10 TFLOPs would significantly decrease inference speed, and $C_{m}=94$, as it is the memory of the Nvidia H100 NVL GPU.
More details about the implementation and setup of the base model and individual PEFT methods are provided in Appendix A.
5 PEFT-Bench Results
Method
# Trainable
Parameters
GLUE (F1/Pearson) SuperGLUE (F1)
@@ 5 PEFT-Bench Results
Prefix Tuning 34,177,536 48.0 12.7 92.1 91.6 82.8 38.8 72.5 45.5 26.9 45.5 72.2 60.1 0 49.8 7.5
P-Tuning 53,130,752 85.7 78.2 81.6 94.7 3.1 89.4 0 85.8 80.4 85.3 86.4 70.8 0 6.0 92.3
[…]
Table 1: Results of PEFT-Bench with different PEFT methods on NLU tasks from GLUE and SuperGLUE benchmarks. We measure Pearson correlation for STS-B and F1 for others.
Method
# Trainable
Parameters
Others Math Code Generation
MMLU PIQA SIQA HS WG OBQA MQA GSM8K SVAMP Conala CAPy APPS
llama-3-8b-instruct

### v2

@@ Abstract
Despite the state-of-the-art performance of Large Language Models (LLMs) achieved on many tasks, their massive scale often leads to high computational and environmental costs, limiting their accessibility. Parameter-Efficient Fine-Tuning (PEFT) methods address this challenge by reducing the number of trainable parameters while maintaining strong downstream performance. Despite the advances in PEFT methods, current evaluations remain limited (in terms of evaluated models and datasets) and difficult to reproduce. To bridge this gap, we introduce PEFT-Bench, a unified end-to-end benchmark for evaluating diverse PEFT methods on autoregressive LLMs. We demonstrate its usage across 27 NLP datasets and 7 PEFT methods. To account for different PEFT training and inference factors, we also introduce the PEFT Soft Cost Penalties (PSCP) metric, which takes trainable parameters, inference speed, and training memory usage into account.
@@ 1 Introduction
Large Language Models (LLMs) (Radford et al., 2019; Achiam et al., 2023; Dubey et al., 2024; Jiang et al., 2023; Team et al., 2024) achieve remarkable results in many Natural Language Understanding (NLU) and Natural Language Generation (NLG) tasks. This significant increase in performance was possible due to the introduction of the attention mechanism forming the transformer architecture (Vaswani, 2017). To achieve such performance, many currently best-performing models contain a vast number of attention blocks, increasing the number of trainable parameters to millions (Devlin et al., 2019), billions (Radford et al., 2019), or even trillions (Fedus et al., 2022). With the increase in parameters, the computational and data requirements for training also increased, making the development of LLMs unavailable for many academic institutions or practitioners, as well as producing a bigger carbon footprint. To address this issue, a new paradigm was introduced, denoted as “pre-train, fine-tune”, in which the language model is at first pre-trained on large amounts of raw data and then, in a fully supervised way, adapted to the downstream task (Liu et al., 2023a). This standard fine-tuning still requires substantial amounts of resources as well as significant quantities of data and storage to store fine-tuned weights for each task.
@@
To ultimately improve and unify the evaluation of PEFT methods, we propose a new benchmark PEFT-Bench [^1] To promote the replicability and usage of our benchmark, we provide a dedicated GitHub repository: https://github.com/kinit-sk/PEFT-Bench., which bridges the gap in the comparison of PEFT methods. We simultaneously introduce the PEFT-Factory framework [^2] PEFT-Factory source code is available at: https://github.com/kinit-sk/PEFT-Factory.  Belanec et al. (2025b), which provides a necessary underlying technological support for execution of the PEFT-Bench benchmark.
@@
  • We introduce the PEFT-Bench, an end-to-end benchmark that defines the datasets, metrics, and methodology for evaluating PEFT methods in NLP in a fair and consistent environment. To the best of our knowledge, we are the first to provide such a benchmark in NLU and NLG tasks for PEFT methods with autoregressive LLMs. To demonstrate its use, we evaluate 7 different PEFT methods in terms of efficiency and stability with limited data on 27 NLP datasets.
@@
  • Finally, we propose the PEFT Soft Cost Penalties (PSCP) metric, which introduces a number of trainable parameters, memory usage, and inference speed in the final score calculation. PSCP reflects the practical feasibility of deploying PEFT methods in real-world scenarios, where efficiency and resource constraints are often as critical as task performance.
@@ 2 Related Work
Many new PEFT methods are primarily evaluated on GLUE and SuperGLUE benchmarks and often include the main results only on non-autoregressive models (Tang et al., 2025; Zhang et al., 2025b, a) (e.g., T5 (Raffel et al., 2020), RoBERTa (Liu et al., 2019)) or include only limited evaluation on autoregressive models Lan et al. (2025); Shi and Lipani (2024); Belanec et al. (2025a). Although some of the new parametrization-based methods (e.g., SVFT (Lingam et al., 2024), SVF (Sun et al., 2025)) are already evaluated on autoregressive models and on natural language generation tasks, a more unified and overarching evaluation is still missing.
@@
Additionally, due to low baseline replicability, we have been witnessing a trend in new soft prompt-based methods to draw numerical results directly from the related works (without actually rerunning the experiments) (Zhang et al., 2025b; Tang et al., 2025). This practice is often not only methodologically incorrect (if the authors do not perfectly match the original environment), but also weakens the results of the proposed method.
@@ 3 PEFT-Bench
Our PEFT-Bench is constructed in multiple steps, which are visualised in the diagram in Figure 1. At a high level, it consists of three main components: 1) a diverse collection of datasets and tasks; 2) representative language models together with a set of PEFT methods; and 3) a set of evaluation metrics.
@@ 3.1.1 Natural Language Understanding and Reasoning Tasks
PEFT-Bench includes 8 datasets from GLUE, split into 5 tasks, namely natural language inference (NLI)MNLI (Williams et al., 2018), QNLI (Rajpurkar et al., 2016), RTE (Dagan et al., 2005; Bar Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009); paraphrase classificationQQP  [^3] https://quoradata.quora.com/First-Quora-Dataset-Release-Question-Pairs, MRPC (Dolan and Brockett, 2005); sentiment classificationSST-2 (Socher et al., 2013); sentence similaritySTS-B (Cer et al., 2017) and acceptability classificationCoLA (Warstadt et al., 2019).
@@
Additionally, PEFT-Bench also includes 7 datasets from SuperGLUE, split into 4 tasks, namely natural language inference (NLI)CB (De Marneffe et al., 2019); question answeringMultiRC (Khashabi et al., 2018), ReCoRD (Zhang et al., 2018), BoolQ (Clark et al., 2019), COPA (Roemmele et al., 2011); word sense disambiguationWiC (Pilehvar and Camacho-Collados, 2019) and coreference resolutionWSC (Levesque et al., 2011);
@@
PEFT-Bench also includes datasets that are, to the best of our knowledge, not part of any well-known benchmarks or are benchmarks on its own, but are still popular and commonly used to evaluate NLU and reasoning in LLMs, totaling 6 datasets split into 3 tasks, namely question answeringMMLU (Hendrycks et al., 2021), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), OBQA (Khot et al., 2019); natural language inference (NLI)HellaSwag (Zellers et al., 2019); commonsense reasoningWinoGrande (Sakaguchi et al., 2021);
@@ 3.1.2 Math Tasks
Another popular task that is used for LLM evaluation is mathematical problem-solving. The difficulty of datasets for such evaluation can range from simple question answering to full mathematical reasoning and problem-solving. PEFT-Bench includes 3 datasets for mathematical problem solving split into 3 tasks, namely question answeringMathQA (Amini et al., 2019); math word problemsGSM8K (Cobbe et al., 2021) and simple math problemsSVAMP (Patel et al., 2021).
@@ 3.2 PEFT Methods and Pretrained Models
With the popularization of PEFT, many new methods are being introduced within a short period of time (Xu et al., 2023; Prottasha et al., 2025). Therefore, it is computationally expensive to evaluate them all. We design our PEFT method selection in PEFT-Bench to cover additive PEFT, reparametrized PEFT, and selective PEFT categories (the categorization of methods that we have selected can be found in Figure 3). To account for each category, we select IA3 (Liu et al., 2022a), Prompt Tuning (Lester et al., 2021), P-Tuning (Liu et al., 2022b), and Prefix Tuning (Li and Liang, 2021) from additive PEFT methods, where IA3 is an adapter-based method, and the rest are soft prompt-based methods. Consequently, we select LoRA (Hu et al., 2022) from reparametrized PEFT methods and LNTuning (Zhao et al., 2024) and BitFit Ben Zaken et al. (2022) from selective PEFT methods. These methods are widely adopted, frequently used as baselines, and together represent one of the most influential and diverse approaches within each PEFT category. We provide a more detailed description of each selected PEFT method in Appendix C.
@@ 3.3 Evaluation metrics
Current works on PEFT methods often measure solely the accuracy performance or the number of trainable parameters, which makes it harder to compare. Moreover, related works only rarely report the memory and inference efficiency in their results. These are the key factors that may affect the preference for using a specific method over another within a certain computational budget. At the same time, for efficient comparison of benchmarked methods, we need to aggregate performance and such additional PEFT training and inference factors into a single metric. To this end, we propose a new PEFT-bencharking metric called PEFT Soft Cost Penalties (PSCP).
We propose to penalize the performance when the costs behind a certain factor are too high. Therefore, we need to scale the performance on the specific task $P_{t}$ with a penalty for that factor, similarly to the Cobb-Douglass production function (Cobb and Douglas, 1928) for $n$ costs:
@@
where $\beta_{i}$ is the importance of $cost_{i}$, and $cost_{i}^{\beta_{i}}$ is the final penalty.
In our case, we consider the following factors: 1) the number of trainable parameters $M_{p}\in\mathbb{N}$; 2) inference speed $M_{f}\in\mathbb{R}^{+}$ (i.e., floating point operations); and 3) average memory usage during training $M_{m}\in\mathbb{R}^{+}$. Since, these factors are not normalized and cannot be used as penalties on their own, we normalize $M_{p}$, $M_{f}$, $M_{m}$ using reference constants $\frac{M_{p}}{C_{p}}$, $\frac{M_{f}}{C_{f}}$, $\frac{M_{m}}{C_{m}}$, where $C_{p}\in\mathbb{N}$, $C_{f}\in\mathbb{R}^{+}$ and $C_{m}\in\mathbb{R}^{+}$. The constants can be set either to a large value that we would not like to achieve (e.g., number of parameters close to full model training, double the inference floating point operations, or maximum memory available for training), or alternatively, they can also be set to the median or geometric mean of all available $M$ values from a specific factor. If we plug this normalization into Equation 1, we would end up with the intermediate equation 2.
@@
This, however, would rescale the performance $P_{t}$ to infinity if $M>>C$, therefore we need to bound it to the interval $[0,1)$ with the reciprocal function (i.e., $x^{-1}$). For this reason, we also need to add 1, so that the result of the reciprocal does not go to infinity when $M<<C$. The final form of our PSCP metric is given by Equation 3.
@@
The proposed metric is easily extensible if additional factors are of interest. In our case, we consider factors that we want to minimize (i.e., the number of parameters, inference speed, and training memory). If we want to include different factors that we want to maximize (i.e., higher is better), we need to switch the numerator and denominator of the fraction for that cost calculation.
As the PSCP allows setting the importance of each cost by setting the $\beta$ value, it makes the PSCP metric further configurable and adaptable for different use cases. If we set the importance of the cost $\beta$ to 0, the resulting penalty will be 1, leaving the final PSCP metric unaffected by that factor. If we set the importance of the cost $\beta$ to a number greater than 0, the resulting penalty for that factor will increase (i.e., $cost_{i}^{\beta_{i}}$ will get closer to 0).
Subsequently, if the importance of all factors is set equally, only the magnitude of the PSCP score changes, not the order of methods. However, this occurs only when the performance differences between the compared methods are high. This can be seen in Appendix B in Table 9. If we want to put more emphasis on a particular cost, we need to set its $\beta$ value greater than that of other costs. This specific behaviour can be seen in Appendix B and Table 10(a), where we decrease the importance of the number of trainable parameters and, therefore, decrease the penalty, which makes LoRA the method with the highest PSCP score.
@@ 3.4 PEFT-Factory
To increase the replicability, modularity, and usability of the PEFT-Bench, we simultaneously introduce PEFT-Factory (Belanec et al., 2025b), a framework for efficient training of autoregressive LLMs, based on one of the popular LLaMA-Factory LLM training tools (Zheng et al., 2024). PEFT-Factory includes new off-the-shelf PEFT methods from the HuggingFace PEFT library (Mangrulkar et al., 2022) and also provides an easy-to-use interface to add newly created PEFT methods to our framework. PEFT-Factory also includes all PEFT-Bench datasets, from which some are specifically adapted for autoregressive generation and classification (e.g., GLUE and SuperGLUE). In addition, PEFT-Factory implements the PSCP metric together with standard metrics and provides automated evaluation and results comparison.
Consequently, PEFT-Factory can be viewed as an instrument of PEFT-Bench that automates the training of different methods in a simple and configurable way. In this manner, PEFT-Factory takes just two configuration files (one for training and one for evaluation) as input, and returns fine-tuned adapter weights alongside the performance metrics. This specific property of PEFT-Factory allows an identical experimental environment during multiple runs (when equivalent hardware configurations are utilized). For this reason, we create a specific training and evaluation configuration for each model and method during the PEFT-Bench execution.
@@ 4 Experimental Setup
Method # Trainable Parameters GLUE (F1/Pearson) SuperGLUE (F1)
@@
Prefix Tuning 34,177,536 48.0 12.7 92.1 91.6 82.8 38.8 72.5 45.5 26.9 45.5 72.2 60.1 0.0 49.8 7.5
@@
[…]
In all datasets, we utilize 10% randomly selected from the training set as a validation set and the original validation set as a test set for evaluation. During the validation phase, we measure validation loss and save the weights of the best validation loss for every 5% of the training steps. We train for 10 epochs with a batch size of 4. We utilize the AdamW (Loshchilov and Hutter, 2019) optimizer with a 0.1 warmup ratio, a weight decay of 1.0e-5, a learning rate of 5.0e-5, and a cosine learning rate scheduler. If not specified, we keep the training hyperparameters equal for all of the experiments and change only the PEFT method hyperparameters.
For IA3, we apply adapters for key, value, and downsample feedforward modules of the transformer architecture. For Prompt Tuning, we fine-tune the soft prompt with a length of 100 tokens and initialize it randomly from the vocabulary of the trained model. For P-Tuning, we also fine-tune the soft prompt with a length of 100 tokens, but we reparametrize the soft prompts with an LSTM model with a hidden size of 768. For Prefix Tuning, we fine-tune the soft prompt with a length of 32 tokens, and we project the soft prompt with a small MLP with a hidden size of 512. For LoRA hyperparameters, we set the rank to 16, alpha to 16, and dropout to 0.05. Similarly to IA3, we also reparametrize the key, value, and downsample matrices of the transformer architecture. For LNTuning, we leave all of the settings to their default values (the modules are selected based on the model architecture). For BitFit, we add the bias terms to the query and value matrices. For the stability experiments, we rerun training for 5 different random seeds, and we train each method for 20 epochs on each dataset.
To calculate the PSCP metric, we calculate $P_{t}$ as an average of the performance metric (depending on the particular dataset) across all 5 runs. Since all performance metrics for each dataset are in the range <0,1><0,1>, we can average the individual performance scores into a single value $P_{t}$ in a fair manner (as is already done in established benchmarks such as GLUE and SuperGLUE). We set $\beta$ for all factors to 1. For the efficiency factors, we set $M_{p}$ as the number of trainable parameters of the PEFT model (counting the parameters that are enabled in gradient descent). Consequently, $M_{f}$ is set as the difference between base model FLOPs and PEFT model FLOPs. Lastly, $M_{m}$ is set as an average maximum memory across all runs.
We set all reference constants $C$ to maximum values that would still be meaningful for training PEFT methods. More specifically, we set the number of trainable parameters $C_{p}=5\times 10^{8}$, as some of the smallest LLMs (e.g., Qwen2.5-0.5B (Yang et al., 2024)) contain 0.5 billion parameters, and if PEFT methods were to train such a number of parameters, it may be more beneficial to perform a full fine-tuning instead. We set the number of floating-point operations during inference $C_{f}=10$ TFLOPs, approximately twice the number of the base model floating-point operations, as going beyond such a number will make inference with PEFT methods unacceptable. Finally, we set the training memory $C_{m}=94$ GB, as it is typically the maximum memory available on high-end GPU cards (e.g., Nvidia H100 NVL GPU). More details about the implementation and individual PEFT methods are provided in Appendix A.
5 PEFT-Bench Results
Method # Trainable Parameters Others Math Code Generation
MMLU PIQA SIQA HS WG OBQA MQA GSM8K SVAMP Conala CAPy APPS
llama-3-8b-instruct
Base N/A 57.1 46.0 70.3 62.2 9.6 78.3 26.8 79.2 56.7 31.2 32.7 21.6
IA3 196,608 64.6 86.0 79.2 91.7 81.4 83.7 44.3 68.3 85.3 23.9 32.0 18.6
[…]
Prefix Tuning 34,177,536 15.4 45.3 32.0 24.9 27.5 50.1 15.5 37.8 85.7 27.5 9.1 15.2